OpenStudy (anonymous):

Which equation is in standard form & represents a line with slope -1 through the point (-3, 5)? A) -x + y = 5 B) y = -x + 5 C) x + y = 2 D) y = -x + 2

OpenStudy (mathstudent55):

A) y = x + 5 (does not have slope -1, can be eliminated) B) y = -x + 5 (has slope -1) C) y = -x + 2 (has slope -1) D) y = -x + 2 (same eq as eq C) )

OpenStudy (mathstudent55):

Test point (-3, 5) in equations B, C, and D. Also, what does your textbook consider standard form for a linear equation?

OpenStudy (tanner23456):

you can use the point slope equation again (y-y1) = m(x-x1) (y-5) = -1(x-(-3)) y-5 = -x-3 slope-intercept form y = -x +2 standard form x+y = 2

OpenStudy (anonymous):

@tanner23456 Okay, so the answer is C because they ask for standard form. Could you help me with a few more??

OpenStudy (tanner23456):

yeah, no worries

OpenStudy (anonymous):

Do you want me to post them here or on another post?

OpenStudy (tanner23456):

I don't care. Whatever you want

OpenStudy (anonymous):

A sapling, or young tree, is planted in a garden. After 3 years, it is 180 cm tall. After 7 years, it is 368 cm tall. How tall will the tree be after 10 years? A) 587 cm B) 509 cm C) 470 cm D) 548 cm

OpenStudy (tanner23456):

you could set up and equation something like this y = mx +b x = number of years y = height first we need to find the slope (y2-y1)/(x2-x1) (368-180)/(7-3) 188/4 m = 47 since the tree's height starts at zero the y-intercept is 0 y= 47x now plug in x y=47(10) y=470

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

An empty 6-gal water jug weighs 0.5 lb. With 3 c of water inside, the jug weighs 2 lb. Which equation models the jug's weight y when it contains x cups of water? A) y = 0.5x - 0.5 B) y = -0.5x - 0.5 C) y = 0.5x + 0.5 D) y = -0.5x + 0.5

OpenStudy (tanner23456):

same type of thing. x= cups of water y= weight find the slope (y2-y1)/(x2-x1) (2-0.5)/(3-0) 1.5/3 slope = 0.5 this time the weight starts at 0.5 when x=0 so the y-intercept will be 0.5 y= 0.5x + 0.5

OpenStudy (anonymous):

What is the slope for the equation y = 7? A) 0 B) -7 C) Undefined D) 7

OpenStudy (tanner23456):

If you think about it on the graph y=7 would look something like this.  therefore the slope is 0